# Diversity or Precision? A Deep Dive into Next Token Prediction

### v1

@@ B.2 Experiment Results
We provide detailed evaluation results to illustrate performance trajectories during the RL process. Tables 14 to 23 display the RL results across different models and training steps.
@@
Table 15: RL performance of the 10B-A0.5B MoE Model with $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.34 0.47 0.34 0.44 0.63 0.57 0.55 0.55 0.65 0.65
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.33 0.00 0.00
Pass@64 12.88 12.94 11.42 11.68 12.93 11.47 9.59 9.17 10.00 10.84
AIME25 Avg@128 0.13 0.23 0.16 0.29 0.36 0.47 0.60 0.44 0.63 0.57
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 6.68 11.27 8.35 14.80 16.58 19.10 18.67 14.10 16.66 16.59
Cons@128 20.00 25.00 17.50 27.50 25.00 30.00 27.50 35.00 32.50 37.50
Pass@64 75.83 76.02 70.67 71.55 73.75 74.33 73.13 73.22 71.65 73.71
Cons@128 14.81 16.00 17.04 20.00 21.04 21.33 21.19 22.52 24.30 24.89
Pass@64 49.58 49.53 52.16 52.06 52.69 52.50 53.48 53.34 52.67 53.31
Cons@128 45.60 48.40 49.80 52.40 54.40 54.40 55.20 56.20 58.20 58.20
Pass@64 82.96 82.71 83.72 83.39 85.93 85.79 85.64 86.01 86.04 85.75
Minerva Avg@128 5.66 5.70 6.42 6.93 7.22 7.84 7.98 8.28 8.36 8.73
Cons@128 9.93 10.66 11.40 12.50 11.40 13.24 12.87 12.50 12.50 13.97
Pass@64 42.39 40.93 41.04 42.75 43.47 42.89 44.35 44.50 40.90 44.90
Avg@128 8.44 9.24 10.14 11.03 11.79 12.52 12.74 13.36 13.97 14.37
Pass@64 45.05 45.57 44.56 46.04 47.56 47.68 47.48 46.72 46.32 47.52
@@
Table 18: RL performance of the 10B-A0.5B MoE Model with $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.29 0.49 0.86 0.49 0.49 0.86 0.86 0.78 0.86 1.15
Cons@128 0.00 0.00 3.33 0.00 0.00 3.33 0.00 3.33 0.00 3.33
Pass@64 9.15 7.93 8.13 10.84 11.68 14.17 14.59 19.19 15.01 21.27
AIME25 Avg@128 0.16 0.29 0.36 0.36 0.63 0.63 0.89 0.70 0.81 1.04
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 8.35 13.78 15.46 7.92 19.81 17.11 19.20 13.12 14.17 17.21
Cons@128 30.00 35.00 37.50 37.50 37.50 32.50 35.00 35.00 32.50 30.00
Pass@64 75.75 73.72 73.54 76.83 74.15 70.35 71.18 74.50 74.13 78.66
Cons@128 13.33 17.78 20.59 20.89 21.93 24.59 24.15 24.00 25.19 26.96
Pass@64 47.16 49.12 52.08 51.30 50.62 52.86 53.15 52.78 54.66 55.99
Cons@128 45.60 48.60 52.60 54.20 54.20 56.20 57.80 59.40 57.00 58.00
Pass@64 81.19 83.88 85.36 84.99 85.75 86.98 88.10 87.39 87.16 86.89
Minerva Avg@128 5.26 6.36 7.45 8.24 9.02 9.89 10.29 10.48 10.14 9.96
Cons@128 10.29 13.24 13.24 13.24 15.07 15.07 15.81 15.81 16.18 16.91
Pass@64 38.70 41.40 43.51 44.46 43.11 45.72 45.94 45.16 45.82 44.46
Avg@128 8.12 10.16 11.96 13.00 14.12 15.13 16.01 16.00 15.71 16.18
Pass@64 43.38 44.97 46.35 46.06 47.52 47.87 48.69 48.69 48.49 50.75
Table 19: RL performance of the 10B-A0.5B MoE Model with $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.42 0.29 0.65 0.70 0.68 0.78 0.68 1.02 0.94 1.04
Cons@128 0.00 0.00 3.33 3.33 0.00 3.33 0.00 3.33 3.33 3.33
Pass@64 18.31 8.32 16.69 19.19 18.77 14.17 12.51 17.53 20.66 16.58
AIME25 Avg@128 0.26 0.16 0.18 0.34 0.26 0.36 0.29 0.26 0.34 0.68
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 14.19 9.17 10.01 14.41 13.77 14.97 12.11 11.06 13.67 19.02
Cons@128 17.50 25.00 22.50 25.00 25.00 30.00 27.50 30.00 32.50 32.50
Pass@64 70.56 72.46 71.53 78.57 76.84 78.90 79.26 74.73 78.60 77.41
Cons@128 14.81 15.56 16.59 19.11 20.44 21.63 20.44 21.78 22.81 25.19
Pass@64 47.23 50.58 51.54 51.79 52.53 51.78 52.20 52.60 53.27 52.94
Cons@128 44.40 47.40 47.80 50.60 52.00 52.80 52.20 55.00 55.60 57.00
Pass@64 81.18 82.20 82.85 83.56 84.79 84.78 85.65 86.00 85.97 86.24
Minerva Avg@128 5.86 5.92 6.83 7.01 7.51 8.13 7.94 8.38 8.80 9.01
Cons@128 11.40 10.66 11.76 11.03 12.13 12.13 12.50 13.97 13.60 16.91
Pass@64 41.33 42.59 43.28 42.82 44.07 44.88 44.56 44.13 46.17 45.37
Avg@128 8.67 9.75 10.68 11.24 12.10 12.94 12.35 13.34 13.87 14.49
Pass@64 45.47 44.22 45.98 48.39 48.46 48.25 47.72 47.68 49.72 49.59
@@
Table 22: RL performance of the 10B-A0.5B MoE Model with $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$.
@@
Table 23: RL performance of the 10B-A0.5B MoE Model with $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$.

### v2

@@ B.2 Experiment Results
We provide detailed evaluation results to illustrate performance trajectories during the RL process. Tables 18 to 32 display the RL results across different models and training steps.
@@
Table 10: Pre-Training performance comparison across different $\beta$ based on 20B-A1B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
General Knowledge MMLU (Acc.) 28.19 35.81 39.35 41.67 29.77 35.09 39.40 41.53 27.33 35.44 38.75 40.78
MMLU-Pro (Acc.) 9.11 11.98 13.45 16.70 10.41 12.15 14.46 15.32 9.40 13.21 13.64 14.94
NaturalQuestions (EM) 8.83 13.57 15.56 18.39 10.00 14.60 16.18 17.15 9.78 14.32 16.01 18.61
TriviaQA (EM) 26.81 37.23 43.54 47.66 27.31 38.02 44.07 46.75 27.37 38.54 43.38 46.17
Average 18.24 24.65 27.98 31.11 19.37 24.97 28.53 30.19 18.47 25.38 27.95 30.13
Commonsense Reasoning Hellaswag (Acc.) 58.17 64.88 67.68 69.05 58.52 65.31 67.88 69.29 58.35 65.43 67.98 69.47
SIQA (Acc.) 42.89 45.75 45.39 47.80 42.07 44.98 46.72 46.16 42.32 44.98 45.70 46.93
PIQA (Acc.) 75.46 77.48 77.58 78.35 74.48 76.77 77.69 77.86 74.59 76.44 77.26 78.18
WinoGrande (Acc.) 55.25 57.85 60.85 62.90 55.09 60.14 59.91 61.19 55.17 57.77 62.35 64.09
OpenBookQA (Acc.) 36.80 37.80 41.00 40.40 37.00 40.20 40.40 41.20 33.20 38.80 42.40 40.20
CommonsenseQA (Acc.) 27.68 50.29 53.81 61.34 31.20 54.22 58.07 58.23 24.73 40.54 54.95 56.84
Average 49.38 55.68 57.72 59.97 49.73 56.94 58.45 58.99 48.06 53.99 58.44 59.28
Knowledge Average 33.81 40.16 42.85 45.54 34.55 40.95 43.49 44.59 33.27 39.69 43.19 44.71
Logic Reasoning ARC-e (Acc.) 65.49 70.03 72.85 72.60 62.84 70.24 70.58 72.98 66.12 69.32 72.39 74.71
ARC-c (Acc.) 34.64 39.93 41.47 43.34 34.04 40.70 44.02 43.00 35.75 40.87 41.64 43.43
BBH (Acc.) 26.69 29.10 31.84 32.50 26.88 30.16 32.07 33.82 26.36 28.81 29.93 30.75
Average 42.27 46.35 48.72 49.48 41.25 47.03 48.89 49.93 42.74 46.33 47.99 49.63
Mathematics GSM8K (Pass@64) 66.43 79.76 83.15 85.75 59.06 76.57 82.05 81.99 63.46 76.58 78.65 84.61
MATH-500 (Pass@64) 47.59 56.75 62.34 63.64 48.07 55.95 61.74 63.97 44.34 54.57 60.00 64.91
Minerva (Pass@64) 17.75 23.32 23.68 25.16 18.77 22.33 22.85 25.82 17.05 19.88 23.71 25.40
OlympiadBench (Pass@64) 25.22 25.88 28.94 27.65 26.30 26.15 28.04 28.86 24.03 25.05 26.47 27.16
Average 39.25 46.43 49.53 50.55 38.05 45.25 48.67 50.16 37.22 44.02 47.21 50.52
Coding HumanEval+ (Pass@64) 26.66 35.66 39.29 40.89 26.66 35.66 39.29 40.89 26.02 32.82 43.16 41.45
MBPP+ (Pass@64) 56.54 71.21 76.10 79.56 60.90 73.75 77.36 78.47 57.42 72.06 76.92 79.34
Average 41.60 53.44 57.70 60.23 43.44 54.60 58.64 59.74 41.72 52.44 60.04 60.40
Reasoning Average 41.04 48.74 51.98 53.42 40.91 48.96 52.07 53.28 40.56 47.60 51.75 53.51
Average 38.15 45.31 48.33 50.27 38.37 45.76 48.63 49.80 37.64 44.43 48.32 49.99
Table 11: Pre-Training performance comparison across different $\tilde{\lambda}$ and $\hat{\lambda}$ based on 20B-A1B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
General Knowledge MMLU (Acc.) 28.84 36.46 39.82 42.04 29.77 35.09 39.40 41.53 25.92 35.71 39.21 41.47
MMLU-Pro (Acc.) 9.77 12.92 15.24 16.20 10.41 12.15 14.46 15.32 9.53 12.26 13.68 15.35
NaturalQuestions (EM) 9.83 14.40 16.18 18.39 10.00 14.60 16.18 17.15 9.06 14.46 15.65 18.95
TriviaQA (EM) 27.44 37.48 42.76 46.19 27.31 38.02 44.07 46.75 26.42 38.38 43.95 47.16
Average 18.97 25.32 28.50 30.71 19.37 24.97 28.53 30.19 17.73 25.20 28.12 30.73
Commonsense Reasoning Hellaswag (Acc.) 57.75 64.03 67.51 69.16 58.18 64.73 68.13 69.10 58.52 65.31 67.88 69.29
SIQA (Acc.) 42.67 46.57 48.62 47.49 42.07 44.98 46.72 46.16 44.22 45.09 47.39 47.70
PIQA (Acc.) 74.86 77.42 77.31 78.24 75.41 77.20 77.09 78.73 74.48 76.77 77.69 77.86
WinoGrande (Acc.) 55.09 58.88 60.06 61.01 55.09 60.14 59.91 61.19 55.41 59.91 61.72 62.59
OpenBookQA (Acc.) 34.80 36.40 41.00 38.60 37.00 40.20 40.40 41.20 36.20 40.00 41.40 42.00
CommonsenseQA (Acc.) 30.22 50.04 56.84 59.71 31.20 54.22 58.07 58.23 22.93 44.55 53.64 58.31
Average 49.23 55.56 58.56 59.04 49.73 56.94 58.45 58.99 48.73 55.25 58.23 59.74
Knowledge Average 34.10 40.44 43.53 44.87 34.55 40.95 43.49 44.59 33.23 40.22 43.18 45.24
Logic Reasoning ARC-e (Acc.) 65.03 69.28 71.72 73.06 62.84 70.24 70.58 72.98 65.82 70.83 70.37 72.22
ARC-c (Acc.) 34.47 39.08 41.47 42.83 34.04 40.70 44.02 43.00 35.84 41.89 41.30 43.09
BBH (Acc.) 26.25 29.15 30.78 33.56 26.88 30.16 32.07 33.82 23.78 29.17 32.07 33.41
Average 41.92 45.84 47.99 49.82 41.25 47.03 48.89 49.93 41.81 47.30 47.91 49.57
Mathematics GSM8K (Pass@64) 61.97 76.10 80.08 84.39 59.06 76.57 82.05 81.99 64.57 76.15 80.86 82.14
MATH-500 (Pass@64) 46.60 56.29 63.10 65.07 48.07 55.95 61.74 63.97 47.49 56.82 61.07 65.56
Minerva (Pass@64) 19.71 20.51 22.84 23.54 18.77 22.33 22.85 25.82 17.63 22.25 23.89 23.90
OlympiadBench (Pass@64) 22.31 24.67 26.50 29.00 26.30 26.15 28.04 28.86 25.74 26.91 26.86 28.54
Average 37.65 44.39 48.13 50.50 38.05 45.25 48.67 50.16 38.86 45.53 48.17 50.04
Coding HumanEval+ (Pass@64) 24.73 34.46 41.83 40.47 25.98 35.44 39.91 41.00 25.83 36.66 40.12 40.79
MBPP+ (Pass@64) 59.21 72.34 76.80 80.47 60.90 73.75 77.36 78.47 58.90 70.87 75.22 77.69
Average 41.97 53.40 59.32 60.47 43.44 54.60 58.64 59.74 42.37 53.77 57.67 59.24
Reasoning Average 40.51 47.88 51.81 53.60 40.91 48.96 52.07 53.28 41.01 48.86 51.25 52.95
Average 37.95 44.90 48.50 50.11 38.37 45.76 48.63 49.80 37.90 45.41 48.02 49.86
@@
Table 16: Mid-Training performance comparison across different $\beta$ based on 20B-A1B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
# Mid-Trained Tokens 25B 50B 75B 100B 25B 50B 75B 100B 25B 50B 75B 100B
General Knowledge MMLU (Acc.) 43.19 44.69 45.71 46.25 43.40 45.56 46.02 46.55 41.97 43.58 44.97 45.42
MMLU-Pro (Acc.) 19.97 23.11 25.02 25.66 21.63 23.73 25.38 25.64 19.34 22.32 24.54 24.48
NaturalQuestions (EM) 16.62 18.06 17.89 18.61 17.09 17.59 17.81 17.92 16.90 16.81 17.42 17.89
TriviaQA (EM) 45.35 46.57 47.31 47.98 45.95 46.65 47.89 48.31 44.76 45.73 47.02 47.10
Average 31.28 33.11 33.98 34.63 32.02 33.38 34.28 34.60 30.74 32.11 33.49 33.72
Commonsense Reasoning Hellaswag (Acc.) 67.88 67.61 68.01 68.61 67.92 68.10 68.45 68.49 67.87 68.12 68.33 68.50
SIQA (Acc.) 44.98 44.68 45.24 44.73 45.70 46.93 46.93 46.88 44.27 46.42 45.50 45.39
PIQA (Acc.) 77.31 76.88 77.80 77.86 77.53 77.31 78.89 78.56 77.80 77.26 77.37 76.77
WinoGrande (Acc.) 61.80 61.56 62.59 63.22 62.19 62.98 63.85 62.59 62.90 64.09 62.90 63.06
OpenBookQA (Acc.) 39.40 39.40 39.60 41.00 41.20 40.80 40.80 40.20 41.20 41.20 40.80 41.00
CommonsenseQA (Acc.) 58.48 59.30 60.52 60.69 61.43 62.33 62.24 61.75 56.67 58.15 59.79 59.87
Average 58.31 58.24 58.96 59.35 59.33 59.74 60.19 59.75 58.45 59.21 59.11 59.10
Knowledge Average 44.80 45.67 46.47 46.99 45.67 46.56 47.23 47.17 44.60 45.66 46.30 46.41
Logic Reasoning ARC-e (Acc.) 70.66 72.22 72.18 73.61 73.02 74.07 74.28 74.12 74.54 74.33 74.20 74.45
ARC-c (Acc.) 43.00 43.86 45.14 45.22 45.48 47.44 47.95 47.35 45.73 46.16 46.93 46.59
BBH (Acc.) 39.99 46.95 49.15 50.25 42.34 46.74 49.90 51.33 39.30 45.65 48.58 49.59
Average 51.22 54.34 55.49 56.36 53.61 56.08 57.38 57.60 53.19 55.38 56.57 56.88
Mathematics GSM8K (Pass@64) 94.64 95.94 95.91 96.43 93.59 95.65 95.69 96.42 94.16 96.04 96.18 96.33
MATH-500 (Pass@64) 77.55 82.14 85.30 85.91 74.76 78.96 80.39 80.26 73.65 78.51 81.04 80.83
Minerva (Pass@64) 31.49 35.35 36.91 37.21 28.43 33.64 35.36 34.61 29.78 32.07 31.38 30.87
OlympiadBench (Pass@64) 38.47 47.15 46.64 47.77 35.99 38.39 38.29 40.91 38.47 45.38 52.48 51.22
Average 60.54 65.15 66.19 66.83 58.19 61.66 62.44 63.05 59.01 63.00 65.27 64.81
Coding HumanEval+ (Pass@64) 67.82 73.38 77.23 79.25 67.09 71.95 76.05 74.99 68.19 71.46 73.26 74.48
MBPP+ (Pass@64) 85.98 88.54 89.77 90.03 87.34 88.88 89.88 90.30 87.75 89.33 89.71 90.32
Average 76.90 80.96 83.50 84.64 77.22 80.42 82.96 82.65 77.97 80.39 81.49 82.40
Reasoning Average 62.89 66.82 68.39 69.28 63.01 66.05 67.59 67.76 63.39 66.26 67.78 68.03
Average 55.65 58.36 59.62 60.36 56.07 58.26 59.45 59.53 55.87 58.02 59.19 59.38
Table 17: Mid-Training performance comparison across different $\tilde{\lambda}$ and $\hat{\lambda}$ based on 20B-A1B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
# Mid-Trained Tokens 25B 50B 75B 100B 25B 50B 75B 100B 25B 50B 75B 100B
General Knowledge MMLU (Acc.) 42.79 44.69 45.86 46.77 43.40 45.56 46.02 46.55 42.56 44.47 46.10 45.61
MMLU-Pro (Acc.) 22.57 24.29 25.91 25.83 21.63 23.73 25.38 25.64 21.57 23.55 26.33 26.15
NaturalQuestions (EM) 16.51 16.54 17.73 18.01 17.09 17.59 17.81 17.92 16.98 18.14 18.53 18.92
TriviaQA (EM) 45.43 45.85 47.29 47.82 45.95 46.65 47.89 48.31 45.97 46.90 47.72 48.35
Average 31.82 32.84 34.20 34.61 32.02 33.38 34.28 34.60 31.77 33.27 34.67 34.76
Commonsense Reasoning Hellaswag (Acc.) 67.67 68.00 68.43 68.46 67.92 68.10 68.45 68.49 68.12 67.73 68.49 68.55
SIQA (Acc.) 44.98 46.01 45.70 46.16 45.70 46.93 46.93 46.88 46.98 47.49 47.70 47.39
PIQA (Acc.) 76.88 77.86 77.86 77.80 77.53 77.31 78.89 78.56 77.37 78.45 78.51 78.73
WinoGrande (Acc.) 62.35 62.51 62.59 61.09 62.19 62.98 63.85 62.59 62.67 62.04 61.25 61.40
OpenBookQA (Acc.) 38.60 39.20 40.00 39.80 41.20 40.80 40.80 40.20 41.60 41.00 42.40 43.00
CommonsenseQA (Acc.) 59.46 59.30 60.52 60.77 61.43 62.33 62.24 61.75 57.08 59.13 58.89 59.95
Average 58.32 58.81 59.18 59.01 59.33 59.74 60.19 59.75 58.97 59.31 59.54 59.84
Knowledge Average 45.07 45.83 46.69 46.81 45.67 46.56 47.23 47.17 45.37 46.29 47.10 47.30
Logic Reasoning ARC-e (Acc.) 73.32 73.78 74.37 74.87 73.02 74.07 74.28 74.12 72.77 74.37 75.34 74.96
ARC-c (Acc.) 45.56 45.65 47.78 47.18 45.48 47.44 47.95 47.35 45.39 46.76 48.29 47.44
BBH (Acc.) 39.46 45.83 48.32 50.24 42.34 46.74 49.90 51.33 42.22 46.57 48.40 49.95
Average 52.78 55.09 56.82 57.43 53.61 56.08 57.38 57.60 53.46 55.90 57.34 57.45
Mathematics GSM8K (Pass@64) 94.18 95.34 96.39 96.51 93.59 95.65 95.69 96.42 93.81 95.16 95.47 95.76
MATH-500 (Pass@64) 76.68 78.17 80.38 80.54 74.76 78.96 80.39 80.26 76.81 78.68 83.05 84.19
Minerva (Pass@64) 29.60 31.41 33.18 37.06 28.43 33.64 35.36 34.61 32.35 34.18 39.24 37.26
OlympiadBench (Pass@64) 37.63 44.08 45.41 48.62 35.99 38.39 38.29 40.91 36.29 46.41 50.51 51.08
Average 59.52 62.25 63.84 65.68 58.19 61.66 62.44 63.05 59.82 63.61 67.07 67.07
Coding HumanEval+ (Pass@64) 66.36 70.07 71.99 71.88 67.09 71.95 76.05 74.99 63.55 69.00 74.08 75.41
MBPP+ (Pass@64) 86.64 88.79 89.46 89.97 87.34 88.88 89.88 90.30 85.21 88.59 89.32 89.22
Average 76.50 79.43 80.73 80.92 77.22 80.42 82.96 82.65 74.38 78.80 81.70 82.32
Reasoning Average 62.93 65.59 67.13 68.01 63.01 66.05 67.59 67.76 62.55 66.10 68.70 68.95
Average 55.79 57.68 58.96 59.53 56.07 58.26 59.45 59.53 55.68 58.18 60.06 60.29
@@
Table 23: RL performance of the 20B-A1B MoE Model with $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.34 0.47 0.34 0.44 0.63 0.57 0.55 0.55 0.65 0.65
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.33 0.00 0.00
Pass@64 12.88 12.94 11.42 11.68 12.93 11.47 9.59 9.17 10.00 10.84
AIME25 Avg@128 0.13 0.23 0.16 0.29 0.36 0.47 0.60 0.44 0.63 0.57
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 6.68 11.27 8.35 14.80 16.58 19.10 18.67 14.10 16.66 16.59
Cons@128 20.00 25.00 17.50 27.50 25.00 30.00 27.50 35.00 32.50 37.50
Pass@64 75.83 76.02 70.67 71.55 73.75 74.33 73.13 73.22 71.65 73.71
Cons@128 14.81 16.00 17.04 20.00 21.04 21.33 21.19 22.52 24.30 24.89
Pass@64 49.58 49.53 52.16 52.06 52.69 52.50 53.48 53.34 52.67 53.31
Cons@128 45.60 48.40 49.80 52.40 54.40 54.40 55.20 56.20 58.20 58.20
Pass@64 82.96 82.71 83.72 83.39 85.93 85.79 85.64 86.01 86.04 85.75
Minerva Avg@128 5.66 5.70 6.42 6.93 7.22 7.84 7.98 8.28 8.36 8.73
Cons@128 9.93 10.66 11.40 12.50 11.40 13.24 12.87 12.50 12.50 13.97
Pass@64 42.39 40.93 41.04 42.75 43.47 42.89 44.35 44.50 40.90 44.90
Avg@128 8.44 9.24 10.14 11.03 11.79 12.52 12.74 13.36 13.97 14.37
Pass@64 45.05 45.57 44.56 46.04 47.56 47.68 47.48 46.72 46.32 47.52
Table 24: RL performance of the 20B-A1B MoE Model with $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.29 0.49 0.86 0.49 0.49 0.86 0.86 0.78 0.86 1.15
Cons@128 0.00 0.00 3.33 0.00 0.00 3.33 0.00 3.33 0.00 3.33
Pass@64 9.15 7.93 8.13 10.84 11.68 14.17 14.59 19.19 15.01 21.27
AIME25 Avg@128 0.16 0.29 0.36 0.36 0.63 0.63 0.89 0.70 0.81 1.04
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 8.35 13.78 15.46 7.92 19.81 17.11 19.20 13.12 14.17 17.21
Cons@128 30.00 35.00 37.50 37.50 37.50 32.50 35.00 35.00 32.50 30.00
Pass@64 75.75 73.72 73.54 76.83 74.15 70.35 71.18 74.50 74.13 78.66
Cons@128 13.33 17.78 20.59 20.89 21.93 24.59 24.15 24.00 25.19 26.96
Pass@64 47.16 49.12 52.08 51.30 50.62 52.86 53.15 52.78 54.66 55.99
Cons@128 45.60 48.60 52.60 54.20 54.20 56.20 57.80 59.40 57.00 58.00
Pass@64 81.19 83.88 85.36 84.99 85.75 86.98 88.10 87.39 87.16 86.89
Minerva Avg@128 5.26 6.36 7.45 8.24 9.02 9.89 10.29 10.48 10.14 9.96
Cons@128 10.29 13.24 13.24 13.24 15.07 15.07 15.81 15.81 16.18 16.91
Pass@64 38.70 41.40 43.51 44.46 43.11 45.72 45.94 45.16 45.82 44.46
Avg@128 8.12 10.16 11.96 13.00 14.12 15.13 16.01 16.00 15.71 16.18
Pass@64 43.38 44.97 46.35 46.06 47.52 47.87 48.69 48.69 48.49 50.75
Table 25: RL performance of the 20B-A1B MoE Model with $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.42 0.29 0.65 0.70 0.68 0.78 0.68 1.02 0.94 1.04
Cons@128 0.00 0.00 3.33 3.33 0.00 3.33 0.00 3.33 3.33 3.33
Pass@64 18.31 8.32 16.69 19.19 18.77 14.17 12.51 17.53 20.66 16.58
AIME25 Avg@128 0.26 0.16 0.18 0.34 0.26 0.36 0.29 0.26 0.34 0.68
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 14.19 9.17 10.01 14.41 13.77 14.97 12.11 11.06 13.67 19.02
Cons@128 17.50 25.00 22.50 25.00 25.00 30.00 27.50 30.00 32.50 32.50
Pass@64 70.56 72.46 71.53 78.57 76.84 78.90 79.26 74.73 78.60 77.41
Cons@128 14.81 15.56 16.59 19.11 20.44 21.63 20.44 21.78 22.81 25.19
Pass@64 47.23 50.58 51.54 51.79 52.53 51.78 52.20 52.60 53.27 52.94
Cons@128 44.40 47.40 47.80 50.60 52.00 52.80 52.20 55.00 55.60 57.00
Pass@64 81.18 82.20 82.85 83.56 84.79 84.78 85.65 86.00 85.97 86.24
Minerva Avg@128 5.86 5.92 6.83 7.01 7.51 8.13 7.94 8.38 8.80 9.01
Cons@128 11.40 10.66 11.76 11.03 12.13 12.13 12.50 13.97 13.60 16.91
Pass@64 41.33 42.59 43.28 42.82 44.07 44.88 44.56 44.13 46.17 45.37
Avg@128 8.67 9.75 10.68 11.24 12.10 12.94 12.35 13.34 13.87 14.49
Pass@64 45.47 44.22 45.98 48.39 48.46 48.25 47.72 47.68 49.72 49.59
Table 26: RL performance of the 20B-A1B MoE Model with $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$.
@@
Table 27: RL performance of the 20B-A1B MoE Model with $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$.